ON AN AIRPLANE THAT IS 2/3 FULL, 20% ARE GIRLS, ONE FOURTH OF PASSENGERS ARE WOMEN, 1/8 OF THE PASSENGERS ARE BOYS AND THERE ARE 68 MEN. HOW MANY SEATS ARE THERE ON THE PLANE?
Answer by stanbon(75887)

ON AN AIRPLANE THAT IS 2/3 FULL, 20% ARE GIRLS, ONE FOURTH OF PASSENGERS ARE WOMEN, 1/8 OF THE PASSENGERS ARE BOYS AND THERE ARE 68 MEN. HOW MANY SEATS ARE THERE ON THE PLANE?
----------------
Let the number of seats on the plane be "x".
Number of occupied seats = (2/3)x
-------------------
Girls = (1/5)(2/3)x = (2/15)x
Women = (1/4)(2/3)x = (1/6)x
Boys = (1/8)(2/3)x = (1/12)x
Men = 68
===================
EQUATION:
(2/15)x + (1/6)x + (1/12)x + 68 = (2/3)x
Multiply thru by 60 to get:
8x + 10x + 5x + 60*68 = 40x
17x = 60*68
x = 60*4 = 240 (Number of seats on the plane)
================
Cheers,
Stan H.